OpenStudy (anonymous):

How many grams of magnesium oxide can be produced when 125 g Mg react with 102 g O2?

OpenStudy (anonymous):

Write down the balanced chemical equation first. 2Mg+O2→2MgO Now find out the number of moles of Mg and the number of moles of Oxygen. Mg: 125/24=5.2083 O2: 102/32=3.1875 Now find out which is the limiting reagent, Mg or O2? Clearly, from the balanced chemical equation, O2 is in excess while Mg is the limiting reagent so it will determine what the amount of product will be. As the ratio of Mg to MgO is 1:1, the number of moles of MgO will be 5.208. Multiply this by the molecular mass of MgO and you get the amount in grams. 5.2083×40=208.3g

OpenStudy (aaronq):

good stuff. you should let them work the answer out themselves, though.

OpenStudy (anonymous):

Sure, I'll keep that in mind next time. :)